OpenStudy (anonymous):

Which of the following ordered pairs represents the solution to the system of equations below? -3x+2y=2 7x-3y=7

OpenStudy (anonymous):

A. (2, 4) B. (4, 7) C. (7, 14) D. (8, 13)

OpenStudy (anonymous):

-3(4)+2(7)=2 7(4)-3(7)=7 -12+14=2 28-21=7 2=2 7=7

OpenStudy (anonymous):

Thats not an option

OpenStudy (anonymous):

B. (4,7)

OpenStudy (anonymous):

sorry, i was plugging in to show you and forgot to write the answer